Circular No. 6397

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 J1 R. H. McNaught, Anglo-Australian Observatory, reports the discovery

by Robert Evans of a comet on a U.K. Schmidt Telescope R film exposed by M. J. Drinkwater. The following observations are available:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
May 10.44268 10^{h}53^{m}  −23^{o}20′  Drinkwater
12.57190 10 52 36.^{s}26 −23 06 23′′.2 16 Jerjen
12.58223 10 52 36.18 −23 06 18.5  _{”}

M. J. Drinkwater (Siding Spring). U.K. Schmidt Telescope. Mid-time of 15-min exposure (comet appears as a short, saturated trail). Tail ∼ 0′.5 long toward the northeast.

H. Jerjen (Siding Spring). 1.0-m reflector + CCD. Measurer McNaught. Comet appears as a wide, parabolic fan with coma diameter 0′.6 and a tail 0′.6 long to the north.

FG SAGITTAE J. A. Mattei, AAVSO, reports that this variable is currently fainter

than it has been at any time since 1894 (cf. IAUC 6393). AAVSO data files indicate that FG Sge started to decline in 1992 Aug. from its average maximum magnitude (over ∼ 25 yr) of 9.2 (cf. IAUC 5604, 5609, 5610); by 1992 Oct., it had reached mag 13.7, and since then it had brightened to mag 10.2 several times (without ever fully reaching its ‘normal’ maximum) and then fading again. Its faintest minimum was mag 14.4 in 1994 Aug. Recent observations indicate that in 1996 Jan. it had reached mag 10.2 and started to fade slowly again by the end of Feb. Recent AAVSO visual observations: Feb. 26.2 UT, 10.5 (G. Krisch, Bockenem, Germany); 27.19, 10.9, (W. Kriebel, Leiblfing-Hailing, Germany); 29.13, 11.9 (P. Maurer, Bad Friedrichshall, Germany); Mar. 9.18, 11.9 (B. Granslo, Nordahl Griegs Vei, Norway); 21.18, 12.4 (A. Diepvens, Balen, Belgium); 28.1, 12.5 (J. Speil, Walbrzych, Poland); Apr. 8.0, 13.9 (L. Szentasko, Budapest, Hungary); 22.13, 14.4: (G. Poyner, Birmingham, England); 28.10, 14.3: (Poyner); May 5.07, [13.7 (Poyner); 7.36, 16.0 (R. Royer, Lakewood, CA; CCD + infrared blocking filter). Mattei notes that there are two stars to the east and very near FG Sge; the star located ∼ 8′′–10′′ east is suspected to be variable, and Royer reported this star at mag 12.3 on May 7.36; the star to the east of this suspected variable was then at mag ∼ 15.0.

1996 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green